OpenStudy (anonymous):

MEDALS WILL BE REWARDED! The molar mass of oxygen gas (O2) is 32.00 g/mol. The molar mass of C3H8 is 44.1 g/mol. What mass of O2, in grams, is required to completely react with 0.025 g C3H8? 0.018 0.034 0.045 0.091

OpenStudy (anonymous):

@whpalmer4 @SolomonZelman Sorry to bother you guys, but I saw that y'all are online and studying chemistry and was wondering if y'all could help me. ?

OpenStudy (solomonzelman):

I suck at science, sorry :(

OpenStudy (anonymous):

It's fine. Thank you for replying though. :)

OpenStudy (whpalmer4):

First, don't we need to know the reaction?

OpenStudy (anonymous):

If I am completely honest, I don't understand this at all.

OpenStudy (whpalmer4):

No problem. We're burning some propane, right?

OpenStudy (anonymous):

I do believe so

OpenStudy (whpalmer4):

We need to make a balanced equation that uses up all of the atoms on the left producing the stuff on the right. nC3H8 + mO2 = C3n + H8n + O2m as a first step

OpenStudy (whpalmer4):

if we let n = 1, so we are reacting 1 propane molecule with an unknown number of oxygen molecules, that gives us C3H8 + mO2 = 3C + 8H + O2m Now the combustion products are going to be what? What do you get when you burn a hydrocarbon like propane in plenty of oxygen?

OpenStudy (whpalmer4):

You get carbon dioxide and water, right? CO2 and H2O

OpenStudy (whpalmer4):

so if we stick with our initial guess of n = 1, we've got 3 carbon atoms to pair up with oxygen. How many oxygen atoms will we need to make the carbon atoms worth of CO2?

OpenStudy (anonymous):

6 atoms of oxygen

OpenStudy (whpalmer4):

Good. That's 3 molecules of O2. C3H8 + 3O2 = 3CO2 + H8 uh, but wait, we've got a balanced equation, but we didn't react all that hydrogen. What are we going to do about that?

OpenStudy (anonymous):

The hydrogen and the other oxygen atoms will combine to make H2O

OpenStudy (whpalmer4):

What other oxygen atoms?

OpenStudy (whpalmer4):

We need to put them in the equation. Can you do so?

OpenStudy (anonymous):

oops sorry I read that wrong, and I'll try.

OpenStudy (anonymous):

Um... C3H8 + 3O2 = 3CO2 +4H2 ?

OpenStudy (whpalmer4):

I thought we were going to make some H2O?

OpenStudy (anonymous):

If we have already used our oxygen atoms to create 3CO2, what oxygen do we have left to combine with the hydrogen to make any H2O?

OpenStudy (anonymous):

@whpalmer4 Hello? Sorry not trying to be pushy.

OpenStudy (whpalmer4):

Well, we need to add some more oxygen to the equation! How much more oxygen do we need to use up 8 H atoms?

OpenStudy (anonymous):

We need 4 oxygen atoms to cancel out all the hydrogen atoms when making H2O

OpenStudy (whpalmer4):

and how many oxygen molecules will that be?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so 2O2

OpenStudy (whpalmer4):

Right. So what is our final equation?

OpenStudy (whpalmer4):

(sorry, my kid is asking me for help with his math homework at the same time, and he naturally gets priority!)

OpenStudy (anonymous):

C3H8 + 5O2 = 3CO2 + 4H20 (and don't worry about it. I know what that's like.)

OpenStudy (whpalmer4):

Yes, that looks good to me. Now for the harder part :-)

OpenStudy (anonymous):

woohoo (sarcasm)

OpenStudy (whpalmer4):

Any questions about how we balanced it, before we go on?

OpenStudy (anonymous):

I think I'm good.

OpenStudy (whpalmer4):

Good, well, feel free to tag me if you get stuck in the future. I'll copy part of the problem here so I don't have to scroll back: The molar mass of oxygen gas (O2) is 32.00 g/mol. The molar mass of C3H8 is 44.1 g/mol. What mass of O2, in grams, is required to completely react with 0.025 g C3H8? How many moles of propane are we reacting?

OpenStudy (anonymous):

Thank you. :) and one mole

OpenStudy (whpalmer4):

Okay, are you sure about that? The molar mass of C3H8 is 44.1 g/mol, which to me suggests that 1 mol of C3H8 weighs in at 44.1 g, but we are only reacting 0.025 g of C3H8...

OpenStudy (anonymous):

Oh yeah...in that case, 1.1025 if I did that right...

OpenStudy (whpalmer4):

Let's think about that for a minute. If I tell you it costs $44.10 to buy a concert ticket, and you check your pockets and tell me that you only have 2 and half cents, do you think you have enough for more than one ticket?!?

OpenStudy (whpalmer4):

\[\frac{0.025\text{ g}}{44.1\text{ g/mol} }= \text{______ mol}\]

OpenStudy (anonymous):

0.00057 I did this one first, but the answer didn't seem right. On the other hand, once you used the ticket analogy, it did.

OpenStudy (whpalmer4):

Yes, that's the right number of moles of C3H8. Now, how many moles of O2 are we going to use?

OpenStudy (whpalmer4):

The coefficient in front of C3H8 is missing, so we assume it is 1, given that we have a balanced equation. The coefficient in front of O2 is 5, so we have 5 moles of O2 for every 1 mole of C3H8.

OpenStudy (anonymous):

5 x 32 = 160 160 x 0.0005668934240362812 (or 0.00057) =0.09070294784580499 (or 0.091) so 0.091

OpenStudy (whpalmer4):

Yes, that's the answer I got. I would have set it up as \[\frac{0.025\text{ g}}{44.1\text{ g/mol}}*5*32\text{ g/mol} = \text{________ g} \]

OpenStudy (anonymous):

Yay! and by the way I realize that I do have a question about balancing the equation: How did you know how many atoms should combine with the hydrogen and with the carbon?

OpenStudy (whpalmer4):

well, we know how many carbon and hydrogen atoms we start with, right?

OpenStudy (anonymous):

8 hydrogen and 3 carbon

OpenStudy (anonymous):

oh so you just fill in as many oxygen atoms as you can with each of the other elements until there are no more O atoms left?

OpenStudy (whpalmer4):

Okay, and we know that we're producing CO2 and H2O, so we do enough O2 to soak up all of the C, and then enough to soak up all of the H

OpenStudy (whpalmer4):

Now sometimes (I can't think of an exact example off the top of my head, unfortunately) you'll end up discovering that you need to add another molecule of the C3H8 (or whatever it is) to get things to balance. That's why I wrote the business with n and m way back when, because I was thinking that this would be such a case. Say we had C3H7 + O2 -> CO2 + H2O if we start the same way, we do 1 C3H7 plus 3 O2 gives us C3H7 + 3O2 -> 3CO2 + 7H Now to use up the 7H, we're kind of stuck, because they get used 2 at a time. So we multiply all the quantities by 2: 2C3H7 + 6O2 -> 6CO2 + 14H now we need 7 O, but that's still not going to work, because O comes as O2, so we double again: 4C3H7 + 12O2 -> 12CO2 + 28H now we can balance that if we add 7O2 to both sides 4C3H7 + 19O2 ->12CO2 + 14H2O left side has 4 *3C = 12C right side has 12C left side has 4*7H = 28H right side has 14*H2 = 28H left side has 19*2O = 38O right side has 12*2O + 14*o = 24+14 O = 38O everything balances

OpenStudy (whpalmer4):

I don't think C3H7 exists, but never mind that :-)

OpenStudy (anonymous):

Um I'm a little confused, but I think I got it. Also, I cannot thank you enough for all the help you've given me. ^_^

OpenStudy (whpalmer4):

You're welcome! I hope some of it sinks in and helps you through the next problem like this. I know how easy it is to think "oh, yeah, that all makes sense" only to be faced with a blank wall as you stare at the sheet of paper trying to figure out how to do it on your own!